OpenStudy (anonymous):

\[evaluate \int\limits\limits \frac{dx}{4x \sqrt{x ^{2}-4^{2}}}\]

OpenStudy (lgbasallote):

you can take out 4 \[\frac{1}{4} \int \frac{dx}{x\sqrt{x^2 -4^2}}\] i believe this is arcsec yes?

OpenStudy (mimi_x3):

trig sub

OpenStudy (anonymous):

yup it is an arcsec.

OpenStudy (lgbasallote):

then problem solved! :D

OpenStudy (mimi_x3):

\(x=4sec\theta\) => \(dx = 4tan\theta sec\theta\)

OpenStudy (anonymous):

the answer must \[\frac{1}{16}\sec ^{-1}\frac{1}{4}x+C\] but how???

OpenStudy (lgbasallote):

sounds right...

OpenStudy (lgbasallote):

\[\int \frac{dx}{x\sqrt{x^2 - 4^2}} \implies \frac{1}{4} \sec^{-1} \frac{1}{4}x\] agree?

OpenStudy (anonymous):

go back to the theorem of \[\int\limits \frac{du}{u \sqrt{u ^{2}-a ^{2}}} ==> \frac{1}{a}\sec ^{-1}\frac{u}{a}+C\]

OpenStudy (anonymous):

@lgbasallote the frac in first term must \[\frac{1}{16}\]

OpenStudy (lgbasallote):

that's because there is still a 1/4 out side

OpenStudy (lgbasallote):

\[\frac{1}{4} [ \frac{1}{4} \sec^{-1} \frac 14 x \implies \frac{1}{16} \sec^{-1} \frac 14 x + C\]

OpenStudy (anonymous):

yah your right. Thanks.

OpenStudy (lgbasallote):

<tips hat>